OpenStudy (jenny.andrade):

Determine the number of solutions the system of 1/2y-5=2 and y=3(2x+5) will have A)no solutions B)impossible to determine C)one solution D)infinite solutions

OpenStudy (zzr0ck3r):

is that \(\frac{1}{2}y\) or \(\frac{1}{2y}\)?

OpenStudy (jenny.andrade):

the first one

OpenStudy (zzr0ck3r):

so you have one equation with a \(\color{blue}{y}\) in it on the left \(\frac{1}{2}\color{blue}{y}-5=2\). Then on the right you have another equation, where \(\color{blue}{y=3(2x+5)}\). So we can take that second equation and plug it into the first one. \(\frac{1}{2}\color{blue}{y}-5=2\) becomes \(\frac{1}{2}3\color{blue}{(2x+5})-5=2\)

OpenStudy (zzr0ck3r):

does this make sense?

OpenStudy (jenny.andrade):

so is it a. b. c or d?

OpenStudy (zzr0ck3r):

can you solve that for x?

OpenStudy (zzr0ck3r):

we are not here to give answers:(

OpenStudy (jenny.andrade):

It doesn't make sense if your trying to solve for x

OpenStudy (zzr0ck3r):

why not?

OpenStudy (jenny.andrade):

srry don't get it

OpenStudy (zzr0ck3r):

\(\frac{1}{2}3(2x+5)-5=2\) distribute the 1/2 \(\\3x+\frac{5}{2}-5=2\) add 5 to both sides \(\\3x+\frac{5}{2}=7\) can you do it from ehre?

OpenStudy (jenny.andrade):

thank you I get it now